# Language Models, Graph Searching, and Supervision Adulteration: When More Supervision is Less and How to Make More More

### v1

@@ Acknowledgments
We thank our supervisor Frank Rudzicz for helping to proofread.
@@
Appendix A Experiments
@@
For each experiment, we report the results of $n=5$ differently seeded trials (except for a few experiments where trials prematurely stopped due to issues with our GPU cluster). We find a high variance for the number of iterations needed to solve the task between trials. [^12] This was also independently observed by Saparov et al. (2025), see Appx. B.2. This makes considering multiple trials important when considering if a given experiment is learnable or not. Note when we say ‘unlearnable’, this does not mean the task is provably unlearnable but rather a shorthand for ‘not found to be empirically learnt given 100 epochs’.
We abuse the term ‘epoch’ to mean 1M samples and do so for reporting results. This is because there are no true epochs when using online datasets. In the original setting using an offline dataset, there are 1M sampled examples, and the models are trained for 100 true epochs.
@@
We implemented our experiments using Fairseq (Ott et al., 2019). We perform greedy decoding via Fairseq’s beamsearch with a single beam. Prefixes up to and including the special start-of-targets, ‘=’ (or start-of-scratchpad, ‘#’) are force-generated. Temperature is set to 1.0 and no length penalty is applied. We generate for max length of the ground-truth sequence plus 20 tokens. Despite common knowledge that beamsearch with a single beam is equivalent to greedy search, we were unsure of this due to specific implementation details about beamsearch, which is complex (Kasai et al., 2024). Both the vanilla and first-come-first-serve variants (with the latter used by Fairseq) should be equivalent with a beamsize of 1 and hence equivalent to greedy search. [^13] As a side note, interestingly, this no longer becomes true if employing the patience hyperparameter ($>1$) proposed in Kasai et al. (2024). This was important to verify as we did not want the model to ‘cheat’ by using post-hoc inference search methods in leu of reasoning.
A.1 Baseline Results
@@
Tbl. 6 provides baseline results of the path-star task (PST) using both edge- and causal-wise shuffling of $G$. We find that the PST is unlearnable, even when using online training, reducing the sample space by setting the vocabulary size to the graph size ($|V|=|G|$), and tokenizing $Q$ before $G$ (Frydenlund, 2024). This is consistent with the results of Bachmann and Nagarajan (2024). Tbl. 7 provides a more fine-grained breakdown of the accuracy of the positional token accuracy for $l_{t}$ and the following two nodes for each run. This shows that $l_{t}$ is predicted at $1/D$ chance, while the next two nodes are predicted with 100% accuracy due to the CHC. This behaviour is illustrated in Fig. 7. One exception is Run 3 of the exp. where $D=5,\,M=7$ which fails to learn the CHC, demonstrating the CHC is not guaranteed to be learnt, despite its seeming simplicity. The start and target nodes are learnt immediately due to positional shortcuts and not because of the CHC.
@@
Table 7: Training positional accuracy for $l_{t}$, pos. 2, and pos. 3 for the edge-wise baseline results in Tbl. 6. $\checkmark$ indicates 100% accuracy. In all but a single run, the CHC is learnt for pos. 2 and 3 (and all other non-leading nodes) and $l_{t}$ is predicted at $1/D$ chance. ‘pos.’ is redacted for space for trials 2-5.
Because the PST in its original form is unlearnable (for reasons not due to regular hyperparamaters), it is impossible to hyperparameter tune it. As such, we also use the causal-wise version to determine valid hyperparameters. The reported results for all experiments are after having found hyperparameters using the causal-wise version of the task. This includes the baseline results which were redone for consistency and to rule-out improper hyperparamaters causing the task to be unlearnable.
A.2 Masking Results
@@
Tbl. 8 provides results using span masking via token dropout, token replacement, and a mixture of both. Mixing performs better than the other two, however, the results are not always consistently better. This is due to the amount of noise being added to the training procedure (and we do not ensure that the same nodes are noised in the same place across the different noise types to control for this). We also show that mixing causal-wise shuffling with masking improves the results over just using either, implying that they are helping to solve different underlying issues (the causal constraint with the former and task decomposition with the latter).
@@
Causal-Wise Token Dropout 5 9 60% 60% 60% 100%
@@
A.3 Alternative Distributions Results
@@
Tbl. 9 provides results using the alternative distributions. We find poor results for LS in particular. We strongly suspect this is because each LS weight also functions as a weight on the corresponding loss term. This means that far-future tokens will have tiny contributions to the overall loss. Although we allow for scaling of the monotonically decreasing terms (via a temperature hyperparameter), we do not experiment with this. It may be that doing so will result in better performance, but, we argue that using RITF instead avoids this complication.
@@
Monotonically Decreasing Label Smoothing (LS) 3 5 100% 100% 100% 100%
@@
Table 9: Alternative (future) distribution results.
A.4 Arm Reconstruction Scratchpads Results
Tbl. 10 shows the results for arm reconstruction scratchpads. These are the first results that see differences between autoregressive inference and teacher-forced inference. Sequence accuracy is evaluated independently for $R_{t}$ and the scratchpad ($S$) in order to support better analysis of what the model is learning. This is why the model can get 100% sequence accuracy for $R_{t}$ but less than that for the SP in the teacher-forced setting and why the autoregressive inference can differ from teacher-forced inference results. As such, ‘Test-Gen $R_{t}$’ is the statistic one should consider when determining if a given experiment actually learnt the task.

### v3

@@ Acknowledgments
We thank Frank Rudzicz and Gagandeep Singh for helping to proofread, and Rich Zemel for suggesting some clarifying experiments.
@@
Appendix A Relationships Between Methods
This section provides a summary of the methods introduced and the various relations between the methods, focusing on how each induces subtask decomposition. We also outline how each method either prevents, alleviates, or just ignores the CHC. Note that, except for the alternative sequential distributions, all methods use standard teacher-forcing and next-token prediction.
A.1 The Different Forms of Masking
(a) The original task with fully observed teacher-forced input induces learning the CHC for all but the leading node. Note, we have extended the length of the arm from Fig. 1.
(b) An example of a fully masked sequence (equivalent to the teacher-less model or NAR in terms of masking). All five predictions are conditioned via different masked tokens (at unique positions). Each prediction must use the backward algorithm to determine the correct target.
(c) A partially masked sequence. Unmasked tokens cause the CHC to be learnt for the corresponding prediction. Predictions after an unmasked token may use the forward algorithm to determine the correct target (indicated with primes).
(d) Alternative sequential distributions create dense decomposition supervision. The CHC is learnt for one prediction at all time-steps, except for the leading node (same as in the original task with no masking). In the first time-step, all predictions corresponding to the fully masked must be learnt; the difference between this and Fig. 12(b) is that all predictions are conditioned from a single hidden-state instead of being separated into five temporally distinct hidden-states created via different masked token inputs. Then, for all other future predictions, the forward algorithm may be used to determine the correct targets since the model conditions on ground-truth tokens on those time-steps. Also, some of these predictions may be repeated at different time-steps. Nodes ‘11’ and ‘2’ are omitted for visibility.
Figure 12: A comparison of subtask decomposition between masking and alternative sequential distributions.
A comparison between fully observed, fully masked, partially masked, and alternative sequential distributions is given in Fig. 12. These are all target-side modifications to the task. Masking modifies the target-side input, while the alternative sequential distributions modify the target-side targets. However, this can also be viewed as inducing the same masking modification on the target-side inputs, except that this is done implicitly instead of explicitly via masked tokens.
We illustrate teacher-forced inputs and corresponding targets while also labeling how the model achieves each prediction in Fig. 12. Predictions learnt by the CHC are indicated with ‘CHC’ red labels. Because the path-star task uses graphs of the same size, the start and target nodes can be trivially predicted via length shortcuts as they are both included in $Q$ (see Fig. 16). These are indicated with ‘LSH’ red labels. Predictions for the second through seventh target tokens that use the desired backward algorithm are indicated with green labels ‘2’ through ‘7’. Those that may use the forward algorithm are indicated with primes and blue labels. Consider again Fig. 5 in Sec. 3.1.2 for a more detailed illustration of how subtask decomposition is achieved via masking.
We illustrate the original adulterated task in Fig. 12(a), where the CHC is learnt for all predictions except for the leading node. The CHC was the mechanism originally attributed to causing the task to be unlearnable by Bachmann and Nagarajan (2024). This motivated the use of a ‘teacher-less’ model, which fully masks the target-side inputs. [^12] This has a technical difference with our masking method. Our masking keeps true autoregressive conditioning at all steps while ‘teacher-less’ models only condition on a single mask token per time-step and, hence, can predict all masked positions in parallel (like a non-autoregressive model). This difference only matters in that our masking method is a standard method to use with language models, while the teacher-less method is not, and this then helps to support our argument that standard methods are sufficient to solve the task and that alternative models are not needed. We illustrate the effect of full masking in Fig. 12(b). This prevents the CHC and induces learning of the desired backward algorithm for all predictions. [^13] Thus, the only methods that completely prevent learning the CHC are full masking, the teacher-less model, and ‘fully’ non-autoregressive models (Frydenlund, 2024). This is in contrast to partial or sampled masking, illustrated in Fig. 12(c), which still allows the CHC to be learnt, hence only alleviating the issue. [^14] This also means that the iterative autoregressive models and the discrete diffusion models only alleviating the issue (see Appx. C.10)., [^15] Bachmann and Nagarajan (2024) introduced the teacher-less model with masking as a direct anti-CHC mechanism on the assumption that it was the mechanism causing the task to be unlearnable and that its prevention is necessary to make the task learnable. However, we show that masking also introduces subtask decomposition. This raises the question: what is the critical aspect of this method that allows for the task to be learnt? We argue that decomposition is the critical aspect, not CHC-prevention. This is held up by the fact that, as mentioned above, partial masking doesn’t actually fully prevent the CHC and still learns the task, and that other methods like alternative sequence distributions, general queries, and general lengths, a), learn the task, b), do not prevent the CHC at all, but, c), do induce decomposition. Hence, decomposition is necessary to learn the task, but preventing the CHC is not, and so, this must be the critical aspect of why masking works.
Whereas full masking induces learning of the backward algorithm for predictions, partial masking means that predictions after any unmasked ground-truth inputs may use the forward algorithm. The differences in algorithms may be important since, while they are very similar, the backward algorithm is required to be learnt to correctly predict the leading node, and as the forward algorithm is different, this means it is not an exact subtask decomposition. Assuming that learning the forward algorithm supports learning the backward algorithm, this suggests that subtask decomposition does not need to be exact mirror of the original or core task. We also discuss this idea in relation to how tree topology induces decomposition in Sec. 3.4. Given the success of these methods, it seems like subtask decomposition does not need to be exact. This makes a formal definition of decomposition difficult (whereas in the exact case, we could just define a task recursively and treat each recursive step as a decomposition).
A.2 Masking and Alternative Sequential Distributions
Comparing masking to alternative sequential distributions described in Sec. 3.2, alternative sequential distributions act like a combination of all three of the described masking types – fully observed (no mask), fully masked, and partially masked – all at once. This is achieved due to having multiple targets per time-step (i.e., being multi-token prediction methods), which leads to a greater density of target supervision (and hence predictions). Or to put another way, more targets are given, and a subset of these targets induce predictions like the fully observed task, a subset induces predictions like the fully masked task, and the remaining subset induces predictions like the partially masked task.
This is illustrated in Fig. 12(d). Here, teacher-forcing with fully observed input is used. This means that the same predictions in Fig. 12(a) are also learnt, thus, the CHC is still learnt. However, by predicting all future tokens from the initial generic task token ‘=’ on the first time-step, no actual instance-specific supervision is being provided to these predictions via teacher-forcing, i.e., there is no conditioning on the correct arm. Thus, these predictions behave exactly like the fully masked predictions in Fig. 12(b). Then, partial ground-truths are observed for future predictions for all other time-steps, and these predictions behave exactly like the partially masked prediction in Fig. 12(c). Note that these are repeated multiple times.
Because alternative sequential distributions act like a combination of masking methods, they induce multihop arm reconstruction and subtask decomposition in the same way as masking. The only difference here is that multiple predictions are conditioned from a single hidden-state instead of being conditioned on five temporally distinct hidden-states (each of which is made unique by using a new masked token and positional embeddings), however, this technical difference does not affect the induced decomposition. This is also why one can view this as implicitly modifying the target-side input with masked tokens (and the galaxy-brain view is that we have converted target-side input supervision to target-side target supervision.).
See Gerontopoulos et al. (2025) for an example of a method which helps bridge the masking method that we use with the alternative sequential distributions to help demonstrate this idea. They introduced extra tokens as input into the model to create new hidden-states from which they can make future predictions. These extra tokens can be compared to masked tokens, except they do not replace exiting tokens. Then this method can be compared to our alternative sequential distributions, where the former creates new hidden-states via modifying the input to the model, whereas the latter reuses hidden-states to make multi-token predictions. This then shows how one can convert input supervision into target supervision using a structured loss. [^16] See Appx. C.11 for a discussion of their work – we highly recommend it to readers who have made it this.
Finally, alternative sequential distributions also create a unique or novel form of subtask decomposition since the same target is reused at different time-steps. For example, it requires a subset of the steps needed to predict a target at step 6 from step 3, compared to predicting the target at step 6 from step 2 (one less step to be precise). Consider the four (teacher-forced) predictions for target ‘63’ indicated with the label $6^{\prime}$ in Fig. 12(d). To predict ‘63’ from ‘59’ requires one hop through ‘1’. To predict ‘63’ from ‘6’ requires two hops; one through ‘59’ and then the same one through ‘1’. To predict ‘63’ from ‘12’ requires three hops; one through ‘6’ and then the same two through ‘1’ and ‘59’ etc.
Because teacher-forcing is still being used for the alternative sequential distributions, the only non-standard aspect of these methods is that they are not strictly using the next-token prediction paradigm but rather a generalization of it – next-tokens prediction, if you will. [^17] This is more commonly called multi-token prediction, but that also includes NAR/IAR/discrete diffusion models that can predict anywhere in the sequence. This distinction is important for this work since the core research question of this work is whether next-token prediction is sufficient to learn the task. Thus, these methods are not directly relevant to that research question. However, they do provide another example of how subtask decomposition makes the task solvable, and this ties back to the core research question, as we show that next-token prediction is sufficient to learn the task, so long as some kind of subtask decomposition is given as supervision and not adulterated. [^18] We also included these methods in this work to explain the success of other multi-token works which consider the path-star task.
A.3 BoW Label Smoothing and Scratchpad
(a) Alternative sequential distribution using uniform label smoothing. This is the same as Fig. 12(d) except the targets are stacked and only prediction types, but not individual predictions, are illustrated for visibility. In each step, ‘2’ and ‘29’ are trivially predicted via the length shortcuts and a token is trivially predicted by the CHC for all but the second time-step. The prediction of the leading node ‘12’ must always be achieved using the forward algorithm.
(b) We sample a BoW order to be used as sequential teacher-forced input into the model. Multiple targets are valid since any unsampled node in the arm can be the next token. Dispite having multiple targets, this is a true next-token distribution.
Figure 13: BoW alternative sequential distribution and an equivalent BoW scratchpad.
As mentioned in Sec. 3.3, the BoW scratchpad unifies the alternative sequential BoW distribution with a single next-token distribution since the next $M$ scratchpad tokens are the BoW. That is, we can treat the multi-target BoW prediction task as a next-token sequential task via intermediate scratchpad predictions. This is done by sampling a permutation of the target arm sequence and treating that as the ground-truth sequence for both the inputs and the targets of the scratchpad.
We apply label smoothing to the scratchpad targets. In theory, this does not need to be done, provided enough BoW orders are sampled. However, given the issue concerning the large sample space of graphs, this would be a bad strategy in practice (especially since the sample space is made even larger when considering sampling a scratchpad in addition to the graph and target arm). By sampling an input order for the scratchpad, we have incorporated order permutation into learning the task. [^19] This relates to encoder-only non- and iterative-autoregressive models, except that these also take it one step further to create true permutation invariance via removing the causal constraint, where this attempts to learn permutation invariance via sampling permuted input sequences but keeps the causal constraint of decoder-only models. Even though the scratchpad has multiple targets, this is not an alternative sequential distribution but a true next-token distribution since having multiple targets is a consequence of the order permutation, which allows for any unsampled node in the arm to be the next token. One can see how this becomes less unified if we weighted the labels like we do with the other alternative sequential distributions, which would result in preferring specific permutations over others.
This is illustrated in Fig. 13, which shows the BoW label smoothing alternative distribution in Fig. 13(a) and how this is equivalent to a BoW scratchpad in Fig. 13(b). Because we allow the BoW scratchpad to be generated in any order, an easy shortcut the model could learn would be to just learn to generate the BoW as the arm in reverse order. We know this shortcut is not learnt since it would allow the model to trivially solve the task (see Tbl. 5).
Now we explain how the BoW scratchpad induces subtask decomposition. First, determining which nodes can be in the BoW requires selecting which nodes belong to the target arm, and this necessitates arm reconstruction. Second, since the trivial reverse order is not learnt, this means that the CHC can not be solely used to achieve this, and that multihop arm reconstruction must be learnt. To see why this is, consider when two nodes, $u$ and $v$ are adjacent in the BoW but not in the arm, then there is no edge, $(u,\,v)$, and to predict $v$ as the next node after $u$ requires going though at least one intermediate node. Then this will induce subtask decomposition in the same way as masking. Fig. 14 illustrates how the next token prediction of the BoW scratchpad is related to masked predictions, where, like the alternative sequence distributions, the masked input is implicit.
Figure 14: On the left, we show a sampled BoW order. This is a permutation of the target sequence, and we show this across all time-steps when using autoregressive next-token prediction. On the right, we show the sampled masks over the ground-truth sequence that are equivalent to a permuted sequence with next-token prediction at each step. Single-target predictions are indicated with red boxes, since this is true of either single- or multi-target predictions.
A.4 Sorting Scratchpad and Graph-Reconstruction Scratchpads
A sorted scratchpad induces subtask decomposition in the same way as the BoW scratchpad as it is also a randomly ordered scratchpad like the BoW scratchpad. This is because the node IDs/values are randomly sampled; therefore, even though the input is in a canonical sorted order, it is still random with respect to the arm’s true order. Thus, masking (including non-autoregressive, iterative-autoregressive, and discrete diffusion models), alternative sequential distributions, the BoW scratchpad, and the sorted scratchpad methods all induce subtask decomposition dispute in the same way, being very different methods in terms of implementation and motivations.
Sorting by itself does not induce subtask decomposition. This is demonstrated with the graph reconstruction scratchpad which learn to either sort the leading or target nodes but fail to learn the task. Here, the scratchpads do not require learning to select which nodes belong in the target arm, and hence there is no decomposition of learning to reconstruct the target arm. Also, the sorting of either the leading or target nodes only works because these can be trivially identified using shortcuts (the leading nodes by their adjacency to the start node and the target nodes by their degree of 1).
Sorting induces node semantics. This is because the model now learns a relationship between all the nodes ($1<2<\dots<|V|$) based solely on their values and not the graph structure. This is in contrast with the other versions of the task, where there is no relationship between nodes outside of the graph structure, making the nodes – and overall graphs – for these versions of the task semanticless.
A.5 Source-Side Methods
So far, the above methods, masking, alternative distributions, and scratchpads, have all been target-side modifications. Trees (or in graph topology in general), general queries, and general lengths [^20] Note that general lengths are actually a graph topology modification as well, just a minor one. are all source-side modifications.
We illustrated how trees induce subtask decomposition in Fig. 8 and how tree topology alleviates the CHC by creating new positions in the sequence which avoid the CHC in Fig. 7. Here, the CHC is not completely prevented. This makes sense since, in the original task, the leading node is immune to the CHC, but all others are not. Then, by inducing subtasks that mirror the original task, we are creating new subtask leading nodes (indicated with primes) that are also immune to the CHC, but any non-leading node in the subtask is still affected by the CHC. The general query and length methods still only have the leading node being immune, thus alleviating the CHC even less.
Figure 15: We show the possible general queries that can be sampled on the right and the corresponding general lengths on the left. Red nodes succeed the given target node and present the only difference between the two versions of the task.
The success of autoregressive language models is due to their sample-efficient training procedure, which use the chain-rule to factorize the probability of a $T$-length sequence into $T$ conditional distributions, which are efficiently trained together in parallel for transformers. One could imagine a poor training procedure for language models, which samples a single conditional distribution to train on per training sample. Like masking, the subtask decomposition of trees is achieved across the sequential dimension, i.e., within a given training sample. This contrasts with general queries and general lengths, which achieve decomposition across training samples, i.e., intra vs. inter sample decomposition. Since the CHC applies to all targets but the leading node, this is effectively sampling a single conditional distribution per training sample – exactly like our imagined poor training procedure. This will make general queries and general lengths less sample efficient than the other methods and may explain their limited success.
We illustrate the possible samples in Fig. 15. General queries and general induce the same possible samples, and hence, the same subtask decomposition. The only difference between them is that the task using general queries also requires the model to generate the full arm with tokens coming after the target token, while general lengths do not. However, this is irrelevant since tokens after the target can trivially be predicted via the CHC.
Appendix B Experiments
@@
We report the results of $n=5$ differently seeded trials for each experiment (except for a few where some trials prematurely stopped due to issues with our GPU cluster). We find a high variance for the number of iterations needed to solve the task between trials. [^21] This was also independently observed by Saparov et al. (2025), see Appx. C.2. This makes considering multiple trials important when considering if a given experiment is learnable or not. Note that when we say ‘unlearnable’, this does not mean the task is provably unlearnable, but rather a shorthand for ‘not found to be empirically learnt given 100 epochs’. We abuse the term ‘epoch’ to mean 1M samples for reporting results as there are no true epochs when using online datasets. In the original setting using an offline dataset, there are 1M sampled examples, and the models are trained for 100 true epochs.
Each trial can be run on a single small GPU (a 12GB Titan X works) and, at max, takes about a day to finish (if the model does not converge and runs for all 100 epochs, and depending on the specific GPU used). The main computational cost is reporting results for 1220 trials. So, while each experiment is cheap and easy to run on moderately sized hardware, the entire set of experiments is not.
@@
We implemented our experiments using Fairseq (Ott et al., 2019). We perform greedy decoding via Fairseq’s beamsearch with a single beam. Prefixes up to and including the special start-of-targets, ‘=’ (or start-of-scratchpad, ‘#’) are force-generated. Temperature is set to 1.0, and no length penalty is applied. We generate for a max length of 20 tokens over the ground-truth sequence length. Despite common knowledge that beamsearch with a single beam is equivalent to greedy search, we were unsure of this due to implementation details about beamsearch, which is complex (Kasai et al., 2024). Both the vanilla and first-come-first-serve (used by Fairseq) variants should be equivalent with a beamsize of 1 and hence equivalent to greedy search. [^22] As a side note, interestingly, this no longer becomes true if employing the patience hyperparameter ($>1$) proposed in Kasai et al. (2024). This was important to verify, as we did not want the model to ‘cheat’ by using post-hoc inference search methods in place of reasoning.
B.1 Baseline Results
Tbl. 1 provides baseline results of the path-star task (PST) using both edge- and causal-wise shuffling of $G$. We find that the PST is unlearnable, even when using online training, reducing the sample space by setting the vocabulary size to the graph size ($|V|=|G|$), and tokenizing $Q$ before $G$ (Frydenlund, 2024). This is consistent with the results of Bachmann and Nagarajan (2024). Tbl. 2 provides a more fine-grained breakdown of the accuracy of the positional token accuracy for $l_{t}$ and the following two nodes for each run. This shows that $l_{t}$ is predicted at $1/D$ chance, while the next two nodes are predicted with 100% accuracy due to the CHC. This behaviour is illustrated in Fig. 16. One exception is Run 3 of the exp. where $D=5,\,M=7$ which fails to learn the CHC, demonstrating the CHC is not guaranteed to be learnt, despite its seeming simplicity. The start and target nodes are learnt immediately due to positional shortcuts and not because of the CHC.
@@ B.1 Baseline Results
Table 2: Training positional accuracy for $l_{t}$, pos. 2, and pos. 3 for the edge-wise baseline results in Tbl. 1. $\checkmark$ indicates 100% accuracy. In all but a single run, the CHC is learnt for pos. 2 and 3 (and all other non-leading nodes) and $l_{t}$ is predicted at $1/D$ chance. ‘pos.’ is redacted for space for trials 2-5.
Because the PST in its original form is unlearnable (for reasons not due to regular hyperparamaters), it is impossible to hyperparameter tune it. As such, we use the causal-wise version to determine valid hyperparameters. The reported results for all experiments are after having found hyperparameters using the causal-wise version of the task and the baseline results which were redone for consistency and to rule-out improper hyperparamaters causing the task to be unlearnable.
B.2 Masking Results
@@
Tbl. 3 provides results using uniform masking via token dropout and span masking via token dropout, token replacement, and a mixture of both. The results show that the difference between these methods are not very consistent. We suspect this is due to the amount of noise being added to the training procedure (and we do not ensure that the same nodes are noised in the same place across the different noise types to control for this). We also show that mixing causal-wise shuffling with masking improves the results over using either in isolation, implying that they are helping to solve different underlying issues (the causal constraint issue with the former and subtask decomposition issue with the latter).
@@
Causal-Wise Span Token Dropout 5 9 60% 60% 60% 100%
@@
B.3 Alternative Distributions Results
@@
Tbl. 4 provides results using the alternative distributions. We find poor results for LS in particular. We strongly suspect this is because each LS weight also functions as a weight on the corresponding loss term. This means that far-future tokens will have tiny contributions to the overall loss. Although we allow for scaling of the monotonically decreasing terms (via a temperature hyperparameter), we do not experiment with this. It may be that doing so will result in better performance, but, we argue that using RITF instead avoids this complication.
@@
Monotonically Decreasing Label Smoothing (LS) 2 5 100% 100% 100% 100%
3 5 100% 100% 100% 100%
4 5 100% 100% 100% 100%
5 5 100% 100% 100% 100%
@@
Table 4: Alternative sequential (future) distribution results.
B.4 Arm Reconstruction Scratchpads Results
Tbl. 5 shows the results for arm reconstruction scratchpads. These are the first results that see differences between autoregressive inference and teacher-forced inference. Sequence accuracy is evaluated independently for $R_{t}$ and the scratchpad ($S$) in order to support better analysis of what the model is learning. This is why the model can get 100% sequence accuracy for $R_{t}$ but less than that for the SP in the teacher-forced setting and why the autoregressive inference can differ from teacher-forced inference results. As such, ‘Test-Gen $R_{t}$’ is the statistic one should consider when determining if a given experiment actually learnt the task.